what is the de Broglie wavelength of a neutron moving at half of the speed of light ? ( the mass of a neutron is 1.67493x10^-27kg)
Answer #1

Speed of light is 3.0*10^8 m/s

so, we have:

v = (3.0*10^8 m/s)/2

= 1.5*10^8 m/s

mass, m = 1.67493*10^-27 Kg

use:

lambda = h / (m*v)

lambda = (6.626*10^-34) / (1.67493*10^-27*1.5*10^8)

lambda = 2.637*10^-15 m

Answer: 2.637*10^-15 m